Progress Report 07/15/03


Home
Up
Emergency Info
Proposals
Facilities
Instrumentation
Systems Status
Schedules
Observer Info
Science at ARO
ARO In The News
Weather
Staff
Education
ARO FAQ
Internal Docs
Tucson Info
Jobs
Photo Gallery
Search ARO
Contacts


Summer Shutdown Progress Report

July 15, 2003

by Employee / Groups.


 

The 10 Meter Gang: (Steve, Bob, Bob, Pat, Mike.)

This is a summary of all that is finished from the summer shutdown list. A similar one should be produced so that we can check off what is/has been done from the Tucson side.

Accomplished:

  1. So far, the biggest accomplishment has been the door and roof seals. The seals between the bottom of the roof on both sides of the building were completed the 2nd of July. Also, the roof seal, running across the top of the right roof, was finished on the 9th of July. These three sets of seals are the only seals that can be done without the aide of a man lift.
  2. Next, the hydraulic cylinder for the left door was rubbing along a lip that spans the bottom of the opening for the cylinder. This lips on both door cylinders were ground off, and seals were installed under the cylinder. This was finished on the 9th of July.
  3. We have almost finish the 'fireproof' the building. We only have to do a few extra little things.The building track has been cleaned. This was accomplished on the 1st of July.
  4. We have inspected and grease the hot-vane slider. It is in good condition. The tracks are in good shape, and the slider seems to run repeatable and reliably.
  5. We changed the oil and checked the o-rings and cleaned the fittings for the for the vacuum pump in the loft. We ran it and checked how well it held a vacuum in the line to the right hand receiver room.
  6. The crane on the fifth floor was opened and examined. There was a 'cooked' bat in the casing of the electronics. Also, we found the maintenance and operation documents. There are a few things that should be done this summer as well as throughout the year to maintain both the door crane and the crane in the loft.
  7. A complete list of the pm's for the 10M is being compiled and recorded for future use.

 

Plan:

  1. The plan for the rest of the week is to start on the doors in the receiver room, clean and grease the AZ and EL gears, retape the insulation on the cassegrain cabin.
  2. Accomplished:
  3. We removed the external charcoal traps for all three compressors in the loft so that they can be brought to town. NOTE: The fittings for the helium lines should only be tightened to 35-45 ft-lbs. The fittings were extremely tight!!!!
  4. Also, discussed with Martin about getting more caps for these lines. They are good for not only keeping dust out but for keeping other contaminates out and also to protect the fittings.
  5. Also, I have been investigating some efficiency measurement on Jupiter with the sis230 receiver at 217, 220, 230, 241, 245 GHz. I have not completed this, but the PSW procedure is giving about 80% at the low end of the band and about 70% at the high end of the band for main beam efficiencies. I noticed also that the BSW would consistently give Tant about 180-200K on Jupiter. This is too high. BVV and I looked into this and found that the calibrations, the sky scan, is lower than the PSW cals. This results in efficiencies with the BSW proc of above 100%, not possible. I will write a memo about this soon. Also, I found that there are two components to the baseline ripple, one at 16.42MHz and one at 11.49MHz. These go with the distance from the cass focus to the subreflector and to the primary respectively. 9347.915mm(second-focus) ~16.046MHz and 12610.8mm(primary- focus) ~11.895MHz. I will investigate this further also, unless anyone has any other suggestions.
  6. BM and SB have been putting the light in the maser chiller room, and ML and BS have been working on greasing the AZ and EL gears. Both of these projects will be done today.
  7. Today, Steve and Bob finished the light in the chiller room and cleaned and serviced the hydraulic system. And, Mike and I finished greasing the AZ and EL gear. That is all.

12 Meter Gang: ( George, Paul, John, Sean, Mike )

  1. Inspected and adjusted all 24 Dome wheels.
  2. Cleaned all ac ducts and filters.
  3. Replaced Onan Batteries.
  4. Changed oil & filters on Onan.
  5. Removed 23mm receiver and sent downtown.
  6. Removed two compressors from telescope.
  7. Removed weight cage.
  8. Inspected and Repaired Dome Drives.
  9. Drained all gear boxes and refilled. ( 10 boxes )
  10. Began work on 480 track.

Kathy Carle:

So far this shutdown I have completed:

  1. Construction, testing, and documenting in AutoCad the new 1-2mm IF Attenuator for new SMT receivers.
  2. Built Fake Telescope test box for D. Forbes use in testing prototype filter bank.
  3. Test and document results for electronic control circuits in 12M Tipper.

The following are being worked on now:

  1. Investigate and requisition replacement lamps for door control panel at SMT. There is still the problem of the red lenses to be solved.
  2. Design, build and test new chopper driver and electronics for 12M.
  3. Design, build and test power supply for CTI frig for 12M.
  4. Build mixer test unit for downtown receiver lab.
  5. Document in AutoCad all new instruments I build. Assemble all data in binders (1 copy for mountain and 1 for downtown.)
  6. Investigate and requisition parts and build other test units when required by engineers rehabilitating 12M receivers and designing new SMT receivers.

William Jonhke:

2MHz Filter banks:

  1. Designed a printed circuit board (PCB) to replace the splitter board in the 2MHz filter banks. Replacement board has been fabricated, stuffed, and preliminary testing has been done. Board requires substitute splitter module from Mini-Circuits. Items have been ordered and now awaiting them to continue. Note: Board was designed to accommodate all the 12M filter banks, but primarily to repair all the 2MHz and the 1MHz filter banks. Working toward having the 2MHz filter banks repaired and aligned by end of summer. Most likely, the 1MHz filter banks as well depending on any additional repairs that they require.

 

Ferrite Modulator Driver:

  1. Prototype test circuit design for Ferrite Modulator Driver for receiver has been completed. Design reviewed with David Forbes. Design appears to be workable. I am currently working on layout of printed circuit board (PCB) for initial prototype. Upon completion for testing with prototype design and identification of required loop gain, a finalized design will be done to fit slot and wiring of card cage.

Tom Folkers:

  1. Began the major job of getting all Linux system up to the same Red Hat release, 7.3, along with the newest kernels. This is no easy task as some systems have special driver needs. Also I'm getting all systems on the "Red Hat Network" which is an semi-automatic package updating system. If any vulnerability is discovered by Red Hat in any of our package, then we are notified and the appropriate new package can be upgraded.
  2. Install a DVD-R drive on smtoast. This replaced the broken CD-ROM and now provides not only CD-ROM capabilities, but CD-R, DVD-R and DVD-ROM. I was able to successfully create and read back a 4.7 Gbytes backup of the ~obs/sys area along with all associated data files. This extends our data export capabilities to DSS-2, DSS-3, CD-R and now DVD-R.
  3. Installed SMT Firewall on Friday July 11. This should afford us increased security of our mountain systems. Over the next few weeks we will be refining our "rules set" to allow normal operation but close out all others. We will also begin to migrate all internal machines to the internal private network. If you are aware of an outside user who can no longer login and use the system, then their machine may need to be added to our database.

Bill Peters:

Summer Shutdown Progress as of 15 July:

  1. The terminal servers no longer need Kronen to download their software upon power up. This is done by Smtocont.
  2. The X-terminals no longer need Kronen to download their software upon power up. This is done by Smtocont.
  3. Hachenburger (the one that had the old Color Monitor display and the Status (SDH)  displays) has been given a flat screen monitor since the original one had the former displays burned into it. It is now showing the TrackerStat display (a duplicate of the display above the Krupp Rack). Waldschloss (the other small one) still has its original monitor and has been configured like Hachenburger. It is available for another (as yet unchosen) display.
  4. The XDM driver has been enabled on Smtoast. Now one can login to Smtoast from Koenigsbacher as smtobs so that a second observer can analyze his data while his colleague uses the Smtoast console.
  5. (Koenigsbacher is the X-terminal with the 21-in display which was used for the Online Chef display and Xephem.)
  6. The bug that caused the "Sambus 80 changed spontaneously" problem and has been squashed. The Sambus bits 24-30 (which include the receiver's Phase Lock Status Bits) are now working. (Pete Hubbard found one of the Sambus units needed repair and replaced it with a spare. And he found that the cable to the Control Room Sambus unit was defective.)
  7. The clearance for the Optical Telescope fine focus stage was found to be sufficient to replace the stage with the model that allows a travel of 1.8 inches. At the next opportunity, the Optical Telescope should be removed so that the current stage (with only 0.5 inch of travel) can be replaced.
  8. The program that logs the UPS status has been ported from Kronen to Smtocont. The porting of the programs that log the wind and the weather is in progress.
  9. I'm ready to zap all comers with the Automatic Defibrillator!

Aldo Apponi:

  1. Following a disastrous 3 months of negations with the NSF to overturn our denial of funding, NSF has recommended that we split up our first proposal into three or four smaller projects and apply to the grants program. Still, this approach does nothing for the outside community where recently we have obtained more than 60 letters supporting our continued operations. Even so, the director of the URO program will not budge on the idea of partial funding of the observatory simply to allow for open access to the community. It is vital that the outside community gain access to our facilities for us to remain competitive in the field.
  2. In response to this situation, I am organizing a proposal which would allow a minimum of 25% open access to the community to conduct science on the KP12m telescope. I have already signed on about 10 co-Is from various institutions in the US including UCBerkeley, UIllinois, UColorado, NRAO, and CalTech to name a few. The proposal will be centered on the uniqueness of the KP12m capabilities, namely frequency coverage in the 65-85 Ghz and 130-180 Ghz with dual channel, single sideband receivers--vitally important to deuterium studies and molecular line searches. Also, highlighted will be the efficient operations (5663 hours on the sky last year with only 4% downtime for weather and 4% downtime for receiver problems), zero-spacing observations for CARMA, remote observing and matching funds from Research Corporation. Although this proposal is somewhat unorthodox for a grants program, it offers some hope to earmark some of the necessary funds to keep the KP12m operating at the 100% level and it will allow for community access where most of our scientific support has come from in the past.

Grahame Gay:

Primary tasks :

  1. Complete the 2 inset, 2mm receiver for Mt Graham
  2. Review the receiver for installation of 1mm high and 1mm low inserts (4)

Steps required for 1 and 2

  1. Replace the existing 1020 refrigerator with a larger capacity balzers.
  2. Design an optics system suitable for 6 inserts with sideband dumps
  3. Complete cross grids and mirror for 2 insert 2mm
  4. Redesign frame for higher optics
  5. Test and evaluate all 6 inserts
  6. Design suitable pump injection system (waveguides and switches) for Gunns.
  7. Interface monitor and control to computer

 

Completed tasks:

  1. Refrigerator has been replaced
  2. Grid design finished and in shop for construction
  3. Optics system partially designed
  4. Interface 90% completed

David Forbes:

I have done the following during summer shutdown:

  1. Talked with Pete Hubbard about replacing AOS IF relays. (Nothing has been actually done yet.)
  2. Installed RedHat 7.3 Linux and RTLinux 3.2 on the filterbank computer smtfb.
  3. Worked a bunch on new filterbanks.
  4. Assisted William J. on old 12M 2MHz filterbank IF splitter replacement card.
  5. Assisted William J. on ferrite modulator driver redesign.

Jon Carlsen:

My main task during summer shutdown thus far has been to assist others in the performance of their duties. I have mainly been tasked with assisting Grahame and Martin with the receivers (namely the new JT system for Mt Graham) and have since taken over the charcoal trap filter regeneration process. Occasionally, I have the opportunity to help out Bill Hale with vehicle issues or running various errands. I look forward to assisting whenever and wherever I am needed to help make summer shutdown a success.


Baltasar Vila-Vilaro:

This is the report of my progress so far.

  1. I have finished moving, re-installation and setting up the right paths for the several software packages needed for the continuum data conversion to the NIC format. They are all now in the /home/smtop/tonic directory. I have also added a couple of new features to the software to make it more robust for single-pixel RXs when used in continuum (for instance, to be able to get a beam map for each of the JT inserts, etc.). Finally, it can also be used (provisionally, see below), to reduce skydips for single pixel RXs.
  2. I have also patched several packages in the GILDAS software frequently reported by observers/operators. Among them, I have now included a beam size calculation for the HHT/SMT in the ASTRO program (this had never worked in any of the releases and we had to use the CSO defaults instead).

TODOS: This is the status of the rest of projects I have in my hands at the moment:

  1. AIPS @ the 12m: I have been tracing the start-up problem that we are experiencing at the 12m. I have checked all the start-up files and seem to be OK. I am now going a bit deeper to see what routine is the one responsible for kicking us out just before the POPS session starts. Hope to have everything clarified by the end of this week.
  2. Solar System Body Ephemeris: I have written the C libraries I will be needing for TT to UT1 conversions, UT1 to JDT, position of body on interpolated osculating elements, etc. At the moment I am about to package the whole thing into a more  compact form and include the new JPL405 ephemeris for the Planets and Minor planets in the final version. We will have to discuss the final interface from tracker.
  3. Skydip reduction of SIS RX data: Finished the basic program structure. At the moment deciding on a user-friendly widget environment for the displaying, etc. The program includes the ATM atmospheric model and will take into account the fact that some of the RXs will be DSB, etc.
  4. (Also writing a paper)

Martin McColl:

  1. The main focus has been on the JT and 1mm receivers and their associated cryogenic systems. The major component has been the change over of Helium refrigerators from the Kitt Peak 1mm to the Mt Graham JT receiver and the associated redesign resulting from these changes.
  2. A charcoal trap baking device has been constructed and regeneration of the two sites charcoal traps has started.
  3. Two Balzers UCH 130's have been rebuilt for the Mt Graham JT and Kitt Peak 2/3mm receivers.
  4. Modifications to the Korn compressor which is to used as an emergency compressor for the CTI 350 systems is also underway.

 Copyright Arizona Radio Observatory.
For problems or questions regarding this web contact [tfolkers{at}email{dot}arizona{dot}edu].
Last updated: 11/08/11.